OpenStudy (anonymous):

SA=2(lh+hw+lw) solve for l

OpenStudy (anonymous):

SA=2(lh+hw+lw) SA=2l(h+w)+2hw l=SA-2hw -------- 2(h+w)

OpenStudy (anonymous):

thank you!